OpenStudy (anonymous):

\[\sqrt{-4} = 2i\] How is this? i = \(\sqrt{-1}\)

OpenStudy (bibby):

That's just how it's defined

OpenStudy (anonymous):

Why does \(\sqrt{4}=2i\)? I don't see how they are equal if 'i' represent sqrt{-1}...

OpenStudy (bibby):

because of the properties of square roots. \[\sqrt-4=\sqrt4*\sqrt-1\]

OpenStudy (anonymous):

Thank you!

OpenStudy (bibby):

No problem